OpenStudy (friendzone):

. Solve: 5x - 6 = 6 + 3x * Remember to isolate the variable!

OpenStudy (immanuelv):

5x - 3x = 6 +6

OpenStudy (immanuelv):

you can handle the rest right @friendzone ?

OpenStudy (friendzone):

@immanuelv you may be a human calculator i dont know nothing because i missed a whole bunch of classes

OpenStudy (immanuelv):

i'm just a human with a calculator lol

OpenStudy (anonymous):

i need one for \(6+6\) too

OpenStudy (friendzone):

@satellite73 yeah

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=6%2B6

OpenStudy (anonymous):

or if you really want to do it with a calculator, then type in the whole thing 5x - 6 = 6 + 3x